Two marbles are drawn without replacement from a box with 3 white, 2 green, 2 red, and 1 blue marble. What is the probability that both marbles are white?
Answer by stanbon(48568)

Two marbles are drawn without replacement from a box with 3 white, 2 green, 2 red, and 1 blue marble. What is the probability that both marbles are white?
Prob(WW without replacement) = (3/8)*(2/7)= 3/28
Cheers,
Stan H.